OpenStudy (anonymous):

Solve for b. 3a + 2b = c b = c - (3a - 2) b = (c - 3a)/2 b = (3a - c)/2

OpenStudy (zzr0ck3r):

3a+2b=c 2b=c-3a now devide by 2 b=c/2-3a/2

OpenStudy (zzr0ck3r):

i dont know how you got c-(3a-2)

OpenStudy (anonymous):

so the second one?

hero (hero):

those are the answer choices lol

OpenStudy (anonymous):

ya?

hero (hero):

1.) 3a + 2b = c 2.) 2b = c - 3a 3.) b = (c - 3a)/2

OpenStudy (zzr0ck3r):

b=(c-3a)/2

Parth (parthkohli):

2nd is the correct option.

Parth (parthkohli):

Can I explain?

hero (hero):

I call myself posting the steps. Step 1 is the original problem. Step 3 is the final result.

OpenStudy (zzr0ck3r):

confused about what your saying

hero (hero):

1.) Original Problem: 3a + 2b = c 2.) Subtract 3a from both sides: 2b = c - 3a 3.) Divide both sides by 2: b = (c - 3a)/2 What more of an explanation do you need?

OpenStudy (zzr0ck3r):

some peole need more explanation i guess, my guess is you did to the first time?

hero (hero):

@zzr0ck3r , I don't recall you posting an explanation at all.

OpenStudy (zzr0ck3r):

I didnt know he needed one, if I did I would not try and make him feel bad about it.

OpenStudy (zzr0ck3r):

I posted the exact thing as you did, but was nice:)

hero (hero):

What you posted was confusing even to me. Anytime you provide an explanation that only confuses further, it is not an explanation. Explanations help people understand the situation more clearly, not confuse them more.

OpenStudy (zzr0ck3r):

and then make fun of them for not knowing? this is your style right>?

OpenStudy (zzr0ck3r):

if what I posested is confusing to you, you might want to study more.

Parth (parthkohli):

@Hero instead of me, Hero should get the medal.

hero (hero):

@zzr0ck3r , I'm not confused by the problem, I'm confused by your attempt at an explanation.

OpenStudy (zzr0ck3r):

I did not explain, I just solved. there is a huge difference.

hero (hero):

Well, next time, I'm sure you'll try to post a more comprehensive and clearer solution.